### v1

# SAP-CoPE: Social-Aware Planning using Cooperative Pose Estimation with Infrastructure Sensor Nodes

Abstract
Autonomous driving systems must operate safely in human-populated indoor environments, where challenges such as limited perception and occlusion sensitivity arise when relying solely on onboard sensors. These factors generate difficulties in the accurate recognition of human intentions and the generation of comfortable, socially aware trajectories. To address these issues, we propose SAP-CoPE, a social-aware planning framework that integrates cooperative infrastructure with a novel 3D human pose estimation method and a model predictive control-based controller. This real-time framework formulates an optimization problem that accounts for uncertainty propagation in the camera projection matrix while ensuring human joint coherence. The proposed method is adaptable to single- or multi-camera configurations and can incorporate sparse LiDAR point-cloud data. To enhance safety and comfort in human environments, we integrate a human personal space field based on human pose into a model predictive controller, enabling the system to navigate while avoiding discomfort zones. Extensive evaluations in both simulated and real-world settings demonstrate the effectiveness of our approach in generating socially aware trajectories for autonomous systems.
Index Terms:
Cooperative Perception, Human Pose Estimation, Social-awareness, Motion Planning, Model Predictive Control (MPC).
I Introduction
Intelligent indoor mobility systems are gaining increasing attention in healthcare, logistics, and the automotive industry [1, 2, 3]. These systems reduce manual transportation tasks, minimizing workplace injuries and improving operational efficiency [4]. Achieving these benefits requires the system to have real-time obstacle detection, human trajectory prediction, and coordinated robot motion for safe and smooth navigation.
While research has enhanced robotic perception ability through high-end onboard sensors, these systems face limitations in perception range and field of view (FOV), especially in crowded indoor spaces. Infrastructure Sensor Nodes (ISNs) offer a solution [5]. Ning et al. proposed ceiling-mounted sensor nodes, which expand the FOV, reduce occlusions, and provide a cost-effective alternative to expensive onboard sensors. A centralized network of ISNs supports multiple robots, improving scalability and reducing costs.
Figure 1: Key challenges: (1) How to reliably detect humans in crowded, occluded scenarios? (2) How to control the robot for human-comfortable motion?
Effective human-robot interaction is essential for both system efficiency and safety [6]. Accurate human intention prediction enables proactive navigation and collision avoidance. To achieve this, We propose a 3D pose estimation approach using probabilistic fusion of image and sparse point cloud data, ensuring real-time, accurate human behavior analysis.
Yang et al. [7] introduced a Model Predictive Control (MPC)-based motion planning algorithm that treats all objects as obstacles, prioritizing only on physical safety. However, psychological safety is also equally important when navigating near humans. Research shows that people maintain social zones in shared spaces [8]. To address this, we integrate a Personal Space concept from psychology into motion planning, enabling robots to respect social behavior, maintain comfortable distances, and move naturally in human-populated environments.
The key contributions of this article are as follows.
  1. 1.
    Real-time Cooperative Perception System: We proposed a cooperative perception framework that leveraging sensor node networks to provide comprehensive environmental awareness. This system accounts for both computation and communication latencies, ensuring a timely, synchronized, and overall understanding of dynamic environments.
  2. 2.
    3D Human Pose Detection We develop a novel approach for 3D human pose detection that combines geometry-based techniques and sensor fusion between image data and sparse point clouds. This method enhances detection accuracy.
  3. 3.
    MPC-based Motion Planning and Control Scheme: We introduced an integrated motion planning and control framework based on MPC. Our approach simultaneously manages both human interactions and obstacle avoidance by incorporating a Personal Space (PS) model for social comfort and Artificial Potential Fields (APF) for dynamic obstacle navigation.
  4. 4.
    Validation through Simulation and Real-world Experiments: The proposed system is extensively verified and validated through simulations and real-world experiments. Results demonstrate its effectiveness in handling human-robot interactions, obstacle avoidance, and motion tracking, enabled by accurate perception.
The structure of this article is organized as follows. Section II presents an overview of the proposed methodology. Section III, details the perception algorithm. Section IV formulates the motion planning and tracking controller. Section V demonstrates the simulation and experimental results generated by the proposed system. Finally, Section VI concludes the study and discusses potential future research directions.
II Related Works
In restricted environments, mobile robots primarily focus on avoiding collisions with objects and are programmed to stop when humans are nearby. However, this strategy is less effective in human-populated areas like hospitals and factories, where interactions with people are unavoidable. This section examines the two key components of these algorithms: perception and planning.
II-A Perception
Cooperative Perception System: Cooperative perception enhances environmental awareness by integrating sensory data from multiple sources, including other robots and infrastructure. This collaborative approach effectively addresses challenges such as occlusions and the limited FOV inherent in individual sensors. Recent advancements highlight the significance of Vehicle-to-Everything (V2X) in autonomous driving domain[9, 10, 11, 12]. Huang et al. provide a comprehensive overview of V2X-based cooperative perception, highlighting its evolution and the critical role of reliable data sharing between robots and infrastructure[10]. However, real-time performance and robustness to communication delays remain challenges. To address these limitations, we propose a real-time cooperative perception system utilizing ceiling-mounted sensors called Infrastructure Sensor Nodes (ISNs). This setup expands the FOV, reduces occlusions in crowded spaces, and ensures timely environmental understanding by mitigating computation and communication latencies, enhancing intelligent indoor mobility systems.
3D Human Pose Detection: Accurately estimating human 3D poses is vital for autonomous systems to interpret and predict human behavior, ensuring safe and socially aware interactions. Single camera-based methods have faced challenges such as depth ambiguity and occlusions[13, 14], thus, the estimation accuracy can be enhanced by leveraging multi-modal sensor data. Zanfir et al. introduced HUM3DIL, a semi-supervised multi-modal approach that embed LiDAR points into pixel-aligned multi-modal features and employing a Transformer architecture for prediction[15]. Similarly, Bauer et al. proposed a weakly supervised method that integrates camera and LiDAR data, enabling accurate 3D human pose estimation without extensive 3D annotations[16], however, it needs dense point cloud to generate high-accuracy pseudo 3D keypoint labels via projecting point cloud to image.
Our work introduces a novel probabilistic 3D human pose estimation method that robustly integrates geometry-based techniques with image data and optionally available sparse point cloud measurements. By leveraging probabilistic fusion and joint coherence. Unlike the previous approaches, our approach reliably provides accurate real-time 3D pose predictions, even in the absence of sparse 3D measurements. This flexibility ensures consistent and precise human behavior understanding, particularly in challenging indoor scenarios characterized by occlusions and low visibility.
II-B Planning
Social-aware Planning: Recent human-aware planning algorithms focus primarily on human interactions while often neglecting static and dynamic obstacles[17, 18, 19, 20]. Among these algorithms, the Social Force Model (SFM) is widely used to generate socially acceptable robot motion. Kamezaki et al. [21] introduced a planning algorithm based on the inducible Social Force Model, while Ratsamme et al. [22] developed the Extended SFM (ESFM) by incorporating the human pose, orientation, etc. to improve the performance of conventional SFM. Although these approaches have shown excellent performance in human-aware planning; however, they do not explicitly incorporate normal obstacles into their planning frameworks. To address this gap, we proposed an MPC-based integrated motion planning and tracking algorithm that simultaneously accounts for both human interactions and obstacle avoidance by leveraging the APF and PS model.
Planning based on Information Types: Human-aware planning algorithms can be broadly categorized into two types: individual state-based information and group-based information. Individual state-based methods use detailed information such as position, pose, orientation, and velocity; for example, Stefanini et al. [23] proposed a novel planning algorithm that incorporates the articulated 3D human poses, semantic labels, and trajectory prediction to describe human behavior. In contrast, group-based methods treat multiple humans as a single entity. Truong et al. [24] opposed a proactive social motion model that accounts socio-spatiotemporal characteristics of human groups, allows the mobile robot to plan a safe trajectory in various situations.
Figure 2: SAP-CoPE Structure Overview
In this paper, we adopt the Personal Space (PS) field proposed by Amaoka et al. [25], which defines a non-viable zone that ensures the psychological safety of humans. This approach incorporates an individual’s position and orientation to construct the PS model into the robot’s motion planning algorithm. While our method is based on individual state information, it achieves performance comparable to group-based approaches, providing a balanced solution for socially aware robot navigation.
III Methodology
The proposed SAP-CoPE method combines real-time cooperative perception with socially aware motion planning and control to enhance autonomous mobility in dynamic indoor environments (Fig. 2). It employs a network of multi-modal ISNs equipped with LiDAR, cameras, and edge computing for local object detection and 3D human pose estimation. A synchronized clock system aligns LiDAR scans with camera operations, ensuring temporally consistent data, which is transmitted to a Central Layer (CL) for global fusion, multi-object tracking, and delay compensation.
The MPC-based controller is designed to generate optimal control commands, enabling the robot to follow a collision-free trajectory while adhering to a predefined reference motion. To enhance navigation, the APF model and PS model are incorporated into the MPC’s objective function. APF ensures effective obstacle avoidance, while the PS model maintains human comfort by respecting personal space boundaries.
The optimal control commands are then transmitted to the Application Layer (AL) for execution. The SAP-CoPE framework is adaptable to various robotic platforms, such as the autonomous cart and autonomous medical bed at the Mechatronic Vehicle Systems Lab, University of Waterloo. In this study, only the autonomous medical bed [26] was used for simulations and experimental validation, and referred as term “robot” through out this paper.
IV Perception
IV-A Local Perception
The local perception consists of camera-based 2D detection, LiDAR point cloud processing, multi-modal class association, and a novel probabilistic 3D human pose estimation framework incorporating joint coherence constraints.
IV-A1 Camera-based 2D Detection
Recent advancements in deep learning have significantly improved camera-based 2D object detection. Notably, YOLOv11 [27] demonstrates superior accuracy and inference speed due to its state-of-the-art backbone and neck architectures. However, the pre-trained model does not generalize well to the top-down sensor node view and is unable to detect the autonomous robot used in our application. To address these limitations, we fine-tuned the YOLOv11 model using a dataset combining COCO [28] and the Indoor Cooperative Infrastructure Dataset (ICID) [5].
IV-A2 LiDAR Processing and Class Association
The LiDAR processing consists of region-of-interest filtering and a hierarchical clustering algorithm to accommodate the scanning pattern. The multi-model class association fuses and aligns the data from both camera and LiDAR modalities to enhance perception accuracy. Further details on the implementation can be found in our previous work [5].
IV-A3 3D Human Pose Estimation with Joint Coherence Constraints
Estimating absolute 3D human pose from a single camera is challenging due to the lack of depth information. In our approach, we first obtain the 2D pixel coordinates of each joint and leverage prior height distribution information to estimate the rough 3D distribution of each joint as $P^{i}\sim N(\mu_{P^{i}},\Sigma_{P^{i}})$ in the absolute world coordinate frame. This estimation is performed using the camera projection matrix $H$, ensuring a more robust and coherent 3D pose representation.
World Coordinate and Jacobian Calculation
Given a pixel point $(x_{p},y_{p})$, height information $z_{w}$, and the camera projection matrix $H$, we aim to estimate the world coordinates $(x_{w},y_{w})$ and the Jacobian matrix $\mathbf{J}$ of $(x_{w},y_{w},z_{w})$ with respect to $(x_{p},y_{p},z_{w})$. The camera projection formula is given by:
$s\begin{bmatrix}x_{p}\\ y_{p}\\ 1\end{bmatrix}=K[R\ t]\begin{bmatrix}x_{w}\\ y_{w}\\ z_{w}\\ 1\end{bmatrix}=H_{3\times 4}\begin{bmatrix}x_{w}\\ y_{w}\\ z_{w}\\ 1\end{bmatrix}$ (1)
where $s$ is a scale factor, $x_{p}$, $y_{p}$ are the pixel coordinates, $K$ is the camera intrinsic matrix, $R$ is the rotation matrix, $t$ is the translation vector, $x_{w}$, $y_{w}$, $z_{w}$ are the world coordinates.
Given a projection matrix $H\in\mathbb{R}^{3\times 4}$, which maps the 3D world coordinates to the 2D image plane, we denote its elements as follows:
$H=\begin{bmatrix}h_{11}&h_{12}&h_{13}&h_{14}\\ h_{21}&h_{22}&h_{23}&h_{24}\\ h_{31}&h_{32}&h_{33}&h_{34}\end{bmatrix}$ (2)
We define the intermediate variables:
$a_{11}$ $=h_{31}x_{p}-h_{11},\quad a_{12}=h_{32}x_{p}-h_{12},$
$a_{21}$ $=h_{31}y_{p}-h_{21},\quad a_{22}=h_{32}y_{p}-h_{22},$
$b_{1}$ $=h_{13}z_{w}-h_{33}x_{p}z_{w}-h_{34}x_{p}+h_{14},$
$b_{2}$ $=h_{23}z_{w}-h_{33}y_{p}z_{w}-h_{34}y_{p}+h_{24}.$
The world coordinates $(x_{w},y_{w})$ are then calculated as:
$x_{w}$ $=\frac{b_{1}a_{22}-b_{2}a_{12}}{a_{11}a_{22}-a_{12}a_{21}}=\frac{N_{x}}{D},$
$y_{w}$ $=\frac{b_{2}a_{11}-b_{1}a_{21}}{a_{11}a_{22}-a_{12}a_{21}}=\frac{N_{y}}{D}.$
The Jacobian matrix $\mathbf{J}$ of $(x_{w},y_{w},z_{w})$ with respect to $(x_{p},y_{p},z_{w})$ is given by:
$\mathbf{J}=\begin{bmatrix}\frac{\partial x_{w}}{\partial x_{p}}&\frac{\partial x_{w}}{\partial y_{p}}&\frac{\partial x_{w}}{\partial z_{w}}\\ \frac{\partial y_{w}}{\partial x_{p}}&\frac{\partial y_{w}}{\partial y_{p}}&\frac{\partial y_{w}}{\partial z_{w}}\\ 0&0&1\end{bmatrix}.$
This Jacobian matrix allows us to propagate uncertainties from the image plane to the world coordinates, essential for precise 3D pose estimation.
Propagation of Uncertainty to 3D World Coordinates
Building upon the computed world coordinates and the derived Jacobian matrix, we propagate the uncertainties of the input parameters $(x_{p},y_{p},z_{w})$ to estimate the 3D distribution of each joint in the absolute world coordinate system.
Using the Jacobian matrix $\mathbf{J}$, the covariance matrix of the estimated 3D position $(x_{w},y_{w},z_{w})$ is obtained through uncertainty propagation as: $\Sigma_{P^{i}}=\mathbf{J}\cdot\Sigma_{(x_{p},y_{p},z_{w})}\cdot\mathbf{J}^{\top}$, and the combined input uncertainty is defined as: $\Sigma_{(x_{p},y_{p},z_{w})}=\mathrm{diag}(\sigma_{x_{p}}^{2},\sigma_{y_{p}}^{2},\sigma_{z_{w}}^{2})$.
The resulting covariance matrix $\Sigma_{P^{i}}$ denotes the uncertainty in the world coordinates of the joint $P^{i}$. Combined with the mean position $\mu_{P^{i}}=[x_{w},y_{w},z_{w}]^{\top}$, we represent the 3D position of each joint as a multivariate Gaussian distribution: $P^{i}\sim\mathcal{N}(\mu_{P^{i}},\Sigma_{P^{i}})$.
Joint Coherence in 3D Pose Estimation
To maintain coherence in the inferred 3D human pose, we impose additional constraints based on the spatial relationships between adjacent joints. This is achieved by incorporating prior knowledge of bone lengths and their associated uncertainties. For example, given two adjacent joints $P^{i}$ and $P^{j}$, the constraint on their Euclidean distance is modeled as: $\|\mu_{P^{i}}-\mu_{P^{j}}\|\approx l_{ij}$, where $l_{ij}$ is the prior mean of the bone length between the joints $i$ and $j$. The uncertainty in $l_{ij}$ is also incorporated, ensuring a consistent and biologically plausible pose.
Final 3D Pose Estimation
By combining the estimated 3D distributions $P^{i}\sim\mathcal{N}(\mu_{P^{i}},\Sigma_{P^{i}})$ of individual joints with joint coherence constraints, we formulate a maximum likelihood estimation (MLE) problem to refine the entire 3D pose:
$\arg\max_{\{\hat{\mu}_{P^{i}}\}}$ $\prod_{i}\exp\left(-\frac{1}{2}(\hat{\mu}_{P^{i}}-\mu_{P^{i}})^{\top}\Sigma_{P^{i}}^{-1}(\hat{\mu}_{P^{i}}-\mu_{P^{i}})\right)$
$\cdot\prod_{i,j}\exp\left(-\frac{\left(\|\hat{\mu}_{P^{i}}-\hat{\mu}_{P^{j}}\|-l_{ij}\right)^{2}}{2\sigma_{ij}^{2}}\right)$
$\cdot\prod_{k\in\mathcal{L}}\exp\left(-\frac{1}{2}(\hat{\mu}_{P^{k}}-\mu_{L^{k}})^{\top}\Sigma_{L^{k}}^{-1}(\hat{\mu}_{P^{k}}-\mu_{L^{k}})\right),$
where the last term represents an additional constraint imposed by LiDAR point cloud measurements. Here, $\mathcal{L}$ denotes the subset of joints for which corresponding LiDAR measurements are available, $\mu_{L^{k}}$ represents the LiDAR-measured 3D joint positions, and $\Sigma_{L^{k}}$ characterizes the measurement uncertainty associated with the LiDAR points.
This comprehensive approach enables robust estimation of absolute 3D human poses from single-camera setups by effectively integrating uncertainties, LiDAR measurements, and enforcing joint coherence constraints.
IV-B Delay-aware Global Perception
The delay-aware global perception framework at the CL is designed to overcome the challenges posed by communication delays in highly dynamic indoor environments. Each sensor node assigns timestamps to detected objects, enabling the system to compensate for delays by comparing these timestamps to the current time at the central node. Using timestamp-based delay compensation, the system adjusts object positions through motion prediction models, ensuring real-time environmental awareness. For human tracking, a non-linear motion model is used to predict movement direction and speed changes. This predictive capability improves the system’s ability to maintain consistent and reliable tracking across different ISNs.
After compensating for delays, the system employs a weighted fusion strategy to combine the local perception results from multiple ISNs. This approach ensures a cohesive and accurate representation of the environment, even when data arrives asynchronously from different sources. By combining advanced local and global perception techniques, the system achieves real-time detection and tracking, effectively compensating for delays while optimizing multi-sensor data fusion in complex indoor environments.
V Integrated Motion Planning and Tracking
Figure 3: Kinematic Motion Model of Robot
V-A Kinematic-based Motion Model
In this paper, a kinematic model is used to describe the robot’s motion as it operates at speeds below 5 km/h. The model is based on the assumptions of tire non-slippage and rigid body dynamics. Notably, the robot is capable of independently controlling the speed and steering angles of both the front and rear wheels, allowing it to have omnidirectional maneuverability. The kinematic model of the proposed robot is formulated as follows.
$\dot{X}=v_{c}\cos(\psi+\beta),\quad\dot{Y}=v_{c}\sin(\psi+\beta)$ (3)
$\dot{\psi}=\frac{v_{f}\sin(\delta_{f})-v_{r}\sin(\delta_{r})}{l_{f}+l_{r}}$ (4)
$\dot{v_{f}}=a_{f},\quad\dot{v_{r}}=a_{r}$ (5)
$\beta=\tan^{-1}(\frac{l_{r}\tan(\delta_{f})+l_{f}\tan(\delta_{r})}{l_{f}+l_{r}})$ (6)
$v_{c}=\frac{v_{f}\cos(\delta_{f})+v_{r}\cos(\delta_{r})}{2\cos(\beta)}$ (7)
Where the $X$, $Y$, and $\psi$ represent robot’s pose in the global frame; $\beta$ is the side-slip angle of C.G., $v_{c}$ is the C.G. velocity; $l_{f}$ and $l_{r}$ are the distances from the C.G. to the front and rear wheels; and $v_{f}$, $v_{r}$, $a_{f}$, $a_{r}$, $\delta_{f}$, and $\delta_{r}$ are the velocity, acceleration, and steering of the front and rear wheels, respectively. In this paper, the system states, $\xi$, system outputs, $\eta$, and system inputs, $u_{c}$ are defined as follows.
$\zeta=\eta=\begin{bmatrix}X&Y&\theta&v_{f}&v_{r}\end{bmatrix}^{T}$ (8)
$u_{c}=\begin{vmatrix}a_{f}&a_{r}&\delta_{f}&\delta_{r}\end{vmatrix}^{T}$ (9)
V-B Personal Space Model
TABLE I: Personal Space Distance Specifications [25]
Types of Distance Values Relationship
Intimate distance 0 - 0.45 m Very intimate relationship
Personal distance 0.45 - 1.2 m Friends
Social distance 1.2 - 3.5 m Strangers
Public distance $>$ 3.5 m Public speaking
This paper adapts the concept of PS[25] to model human behavior in interaction with autonomous systems. Personal space captures an individual’s non-verbal and non-contact communication channels, influencing how they perceive proximity to others. Based on the studies from Edward T. Hall [29] and R. Sommer [30], PS can be divided into four regions: public distance, social distance, personal distance, and intimate distance. These regions define the various zones of comfort and social interaction. The graphical representation and the relative distances can be found in Fig. 4 and Table I.
Figure 4: Graphical Representation of Personal Field. a) Definition of Personal Field in psychology [25]. b) Constructed Personal Field using Mathematical Expression
Let person $P$ is at position $\{x_{p},y_{p}\}$ in the global frame with heading angle $\theta_{p}$; and let another person $Q$ is at position $\{x_{q},y_{q}\}$ in the global frame. The PS model of person $P$ relative to person $Q$ is denoted as $\Omega$, which is divided into two parts, where $\Omega_{f}$ represents the front PS, and the $\Omega_{r}$ is the rear PS. Additionally, $\gamma(x_{p})=1$ if $x_{p}\geqslant 0$ and $\gamma(x_{p})=0$ otherwise. The expression of PS model can be found below.
$\Omega=\gamma(x_{p})\Omega_{f}+(1-\gamma(x_{p}))\Omega_{r}$ (10)
Each front and rear personal space $\Omega_{i},i\in\{f,r\}$ is defined as a 2D Gaussian function using the covariance matrix $\Sigma$, relative distance $d$, and heading angle $\theta$.
$\Omega_{i}=e^{-0.5d^{T}\Sigma_{i}^{-1}d},i\in\{f,r\}$ (11)
$d=\begin{bmatrix}x_{p}-x_{q}\\ y_{p}-y_{q}\end{bmatrix}\begin{bmatrix}cos(\theta_{p})&sin(\theta_{p})\\ -sin(\theta_{p})&cos(\theta_{p})\end{bmatrix}$ (12)
$\Sigma$ defines the shape of the PS. Based on the studies of Shozo[31], it is suggested that the front PS of a human should be twice as large as the rear, left, and right sides. To implement this, the $\Sigma_{f}$ and $\Sigma_{r}$ are defined as follows, where $\sigma_{xx}=2\sigma_{yy}$.
$\Sigma_{f}=\begin{bmatrix}\sigma_{xx}&0\\ 0&\sigma_{yy}\end{bmatrix},\quad\Sigma_{r}=\begin{bmatrix}\sigma_{yy}&0\\ 0&\sigma_{yy}\end{bmatrix}$ (13)
Amaoka et al. [25] stated that the values of $\sigma_{xx}$ should depend on several factors such as age, gender, etc.; however, since this paper is focused on the overall framework for indoor mobility system, the values of $\sigma_{xx}$ is assumed to be fixed at $0.5$ for all people when conducting simulations and experiments.
V-C Model Predictive Controller Design
The formulation of the proposed integrated controller can be expressed using the following equations. Equ. 14 is the cost function of the MPC; Specifically, $J_{1}$ is the APF function for non-human-type obstacles, with the detailed expression provided in our previous work [7, 32]; $J_{2}$ is the PS model for humans; $J_{3}$ and $J_{4}$ are the tracking and control costs, respectively, which are the same as in regular MPC.
Furthermore, Equ. 18 and 19 describe the motion model of the robot; Equ. 20 defines the control rate constraint; Equ. 21 presents the control constraints; Equ. 22 specifies the state constraints; and lastly, Equ. 23 defines the wheel speed constraint to prevent wheel slippage caused by speed differences between the front and rear wheels. More details can be found in [7].
The Taylor series expansion approximates the PS and APF functions into their quadratic form, where the nominal points are the predicted position of the robot, human, and obstacles using the Constant Velocity and Turning Rate (CVTR) over $N_{p}$ time steps. The optimization problem is solved using OSQP[33].
$\arg\min_{u_{c}}J=J_{1}+J_{2}+J_{3}+J_{4}$ (14)
where,
$J_{1}=\sum_{p=1}^{N_{obs}}APF(p),J_{2}=\sum_{q=1}^{N_{human}}PS(q)$ (15)
$J_{3}=\sum_{i=0}^{N_{p}-1}\|\eta(k+i+1)-\eta_{\text{Ref}}(k+i+1)\|_{Q}^{2}$ (16)
$J_{4}=\sum_{j=0}^{N_{c}-1}\|\Delta u_{c}(k+j)\|^{2}_{R}$ (17)
s.t.
$\zeta(k+i+1)=\tilde{A}_{k}\zeta(k+i)+\tilde{B}_{k}\Delta u_{c}(k+j)+\tilde{d}_{k}(k+i)$ (18)
$\eta(k+i+1)=\tilde{C}_{k}\zeta(k+i+1)$ (19)
$\Delta u_{c\min}\leq\Delta u_{c}(k+j)\leq\Delta u_{c\max}$ (20)
$u_{c\min}\leq u_{c}(k+j)\leq u_{c\max}$ (21)
$\eta_{\min}\leq\eta(k+i)\leq\eta_{\max}$ (22)
$\left|E_{K}\Delta u_{c}(k+j)+g(\zeta_{0}(k),u_{0}(k))\right|\leq 0.1$ (23)
$\zeta(k)=\zeta(0)$ (24)
VI Simulation
The parameters for the robot, perception system, and integrated controller are summarized in the table below. TThe proposed perception system and controller were implemented and tested in both simulation and experiments using the Python platform.
TABLE II: Robot and MPC Tuning Parameters
Symbol Value Symbol Value
$l_{f},l_{r}$ 1.2 m $N_{p}$ 20
$u_{c\max}$ $\left[1,1,\frac{\pi}{2},\frac{\pi}{2}\right]$ $N_{c}$ 10
$u_{c\min}$$\left[1,1,\frac{\pi}{2},\frac{\pi}{2}\right]$ $Q$ $diag(2,2,6,10,10)$
$\Delta u_{c\max}$ $\left[0.8,0.8,\frac{\pi}{12},\frac{\pi}{12}\right]$ $R$ $100~{}diag(3,3,4,4)$
$\Delta u_{c\min}$$\left[0.8,0.8,\frac{\pi}{12},\frac{\pi}{12}\right]$ $T_{s}$ 0.1 sec
$\eta_{\max}$ $\left[\text{Inf},\text{Inf},\text{Inf},1.0,1.0\right]$
$\eta_{\min}$$\left[\text{Inf},\text{Inf},\text{Inf},0.1,0.1\right]$
Figure 5: Simulation Results Comparison between Baseline and Proposed Method
Fig. 5 demonstrates the results of the baseline algorithm, which models all obstacles using the APF, and the proposed SAP-CoPE approach. In this simulation, the robot encounters two obstacles: a static obstacle represented as a green rectangle and a human represented as a green circle with the black line indicating their heading angle. The robot’s task is to follow the reference path, defined by the gray dashed line, while avoiding the obstacles.
In the figure, the baseline algorithm successfully avoids the collision with both obstacles while staying close to the reference path, which can be considered a successful outcome; however, when encountering the human-type obstacle, the robot enters the human’s PS after making the 90-degree turn. This may disturb the human and introduce psychological pressure, as well as increase the risk of collision.
In contrast, the SAP-CoPE, which incorporates the personal space model, can avoid collisions and maintain an appropriate distance from humans. It is worth noting that the robot follows the same trajectory when encountering a stationary obstacle as the baseline algorithm, with deviations occurring only when encountering human-type obstacles. This demonstrates the effectiveness of the presented PS model in maintaining human comfort.
VII Experiments
VII-A Perception Experiments
VII-A1 Evaluation Metrics
To rigorously evaluate the performance of the proposed 3D human pose estimation method, we use two principal metrics, compared against annotated ground truth data: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE). The MAE reflects the average magnitude of estimation errors, while RMSE captures the standard deviation and sensitivity to outliers. The evaluation metrics specifically include position error (measured in meters) and yaw error (measured in degrees for orientation).
VII-A2 Results
We compare two pose estimation methodologies: Camera Pose Estimation uses only image-based detections, and Camera-LiDAR Pose Estimation incorporates matched LiDAR points for each joint, if available. Quantitative results are summarized in Table III, while spatial distributions are provided in Figs. 6-7.
TABLE III: Pose Estimation Performance Evaluation
Method Position Error (m) Yaw Error (deg)
MAE / RMSE MAE / RMSE
Camera Pose 0.1166 / 0.1308 8.4912 / 10.2322
Camera-Lidar Pose 0.0555 / 0.0640 8.3785 / 10.0688
VII-A3 Error Analysis
Position Error
The integration of LiDAR data significantly reduces position estimation errors, as measured by MAE and RMSE, by 52.40% and 51.07% Specifically, the Camera-LiDAR method (MAE = 0.0555m, RMSE = 0.0640m) outperforms the camera-only approach (MAE = 0.1166m, RMSE = 0.1308m). This improvement is evident in the spatial distribution of errors (Fig. 6), where the Camera-LiDAR method demonstrates consistently lower positional errors throughout the evaluated region. Nonetheless, the camera-only method still achieves noteworthy performance, underscoring its effectiveness even without LiDAR data.
Yaw Error
Yaw errors are relatively consistent across both methodologies, with the Camera-LiDAR fusion method showing only marginally lower errors (MAE = 8.3785) compared to the camera-only method (MAE = 8.4912). This suggests that orientation estimation benefits minimally from the addition of LiDAR data, in contrast to the substantial improvements observed in positional accuracy. Despite this minor difference, the camera-only yaw estimation remains robust and effective for practical use.
Figure 6: Spatial distribution of position errors. (a) Camera-only pose estimation. (b) Camera-Lidar fusion.
Figure 7: Spatial distribution of yaw errors. (a) Camera-only pose estimation. (b) Camera-Lidar fusion.
VII-A4 Discussion
The results indicate that integrating LiDAR data significantly improves position accuracy but has a limited effect on yaw estimation. This is likely because LiDAR provides precise depth information, which aidis in localization, whereas yaw estimation largely depends on 2D visual cues.
VII-B Overall Experiments
The proposed SAP-CoPE framework was evaluated through three scenarios: obstacle and human interaction, yielding, and group interaction. Figure 8 illustrates an example of the experimental setup, including RViz visualization and left and right images from the ISN. A detailed visualization, including simulation results and all experiments, is available at https://github.com/HopeYless/SAP-CoPE-Project.
Figure 8: Instance in a real-life experiment on the robot, showing the RViz visualization and images captured by an ISN
VII-B1 Scenario One: Obstacle and Human Interaction
Figure 9: Experiment Results of Obstacle and Human Interaction Scenario
In this scenario, the robot encounters two obstacles obstructing its reference path. The first obstacle, representing a stationary object, is positioned at the top right as a green rectangle. The second obstacle is a human, moving from left to right along the scene, represented by a green circle, with the black line indicating their heading angle. The robot’s reference trajectory is defined as the centerline of the hallway, with a reference heading angle of constant 180 degrees. The reference speed for both wheels is set to a constant 0.8 m/s.
In Fig. 9 (a), the solid black markers (circles, triangles, diamonds, and squares) represent different time stamps of the robot and the human. The robot is shown in a blue rectangle. The plot demonstrates that the robot successfully avoided both obstacles, maintaining minimal deviation from its reference trajectory. Fig. 9 (b) and (c) present the robot’s heading angle and wheel linear velocity, respectively. The heading angle plot confirms that the robot consistently tracked the reference signal throughout the experiment, while the velocity plot shows that the robot follows to the reference wheel velocity.
Figs. 9 (d) and (e) demonstrated the optimal acceleration and steering angle generated by the proposed MPC controller. All values remained within the defined constraints, confirming the effectiveness of the control strategy.
VII-B2 Scenario Two: Yielding
Figure 10: Experiment Results of Yielding Scenario
In the second scenario, the planning algorithm was tested to determine whether the robot could appropriately yield to other humans and reduce speed when there was not enough clearance to pass. This scenario involved two humans: one stationary, facing to the right, and another approaching the robot rapidly from behind. The reference trajectory was defined the same as in the previous case.
Fig. 10 (a) demonstrates the robot’s ability to accurately identify the situation. Initially, the robot moved downward to yield to the fast-approaching individual. At the 18th second, both the stationary and dynamic individuals occupied positions that left not enough space for the robot to proceed. In response, the robot correctly decided to reduce speed at an appropriate distance in front of the stationary individual and wait for the dynamic person to pass. Once the space cleared, the robot continued to track its defined reference motion.
Fig. 10 (b) and (c) illustrate the heading angle and wheel linear velocity of the robot, respectively. Both plots show the robot closely following the defined reference signal. Lastly, Fig. 10 (d) and (e) show the optimal acceleration and steering angle generated by the proposed MPC, all of which fall within defined constraints.
VII-B3 Scenario Three: Group Interaction
In the third scenario, the planning algorithm was tested to verify if the algorithm is able to identify a group of people and output appropriate actions. In this scenario, two people were initially stationary and facing each other, leaving sufficient clearance for the robot to pass. The reference trajectory was defined the same as in the previous cases.
Fig. 11 (a) demonstrates the robot’s ability to identify the group of people. Initially, the robot came to a complete stop in front of the two people, even though the clearance between them was enough for the robot to pass. However, the robot recognized that passing between the two individuals could disturb both of them. Once the person in front had walked away, the robot resumed its operation and followed the defined reference trajectory.
Fig. 11 (b) and (c) illustrate the heading angle and wheel linear velocity of the robot, respectively. Both plots show the robot closely following the defined reference signal. Note, in Fig. 11 (c), the plot demonstrates the robot comes to a complete stop around the 7th second. Lastly, Fig. 11 (d) and (e) show the optimal acceleration and steering angle generated by the proposed MPC, all of which fall within defined constraints.
Figure 11: Experiment Results of Group Interaction Scenario
VIII Conclusions and Future Work
In this paper, we propose a SAP-CoPE framework for autonomous driving systems operating in human-populated environments. By leveraging a cooperative infrastructure sensor network and a novel 3D human pose estimation method, our approach effectively addresses perception limitations and occlusion challenges of accurate human intention recognition and safe navigation. Additionally, the integration of human-pose-based PS concepts into an MPC controller enables the generation of trajectories that prioritize both safety and human comfort. Extensive evaluations in both simulation and real-world environments validate the effectiveness of our proposed SAP-CoPE framework by producing socially aware trajectories. Future work will focus on enhancing system robustness in highly dynamic environments and further improving human intention prediction with the combination of learning-based algorithms.
Acknowledgment
The authors gratefully acknowledge the financial support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and MITACS, as well as the financial and technical support provided by Rogers Communications Inc. Canada.